A roof rises 7.5 ft over a horizontal distance of 12.27 ft. What is the slope of the roof to the nearest hundredth?
thank you for your help

Answer by checkley75(3666)

WHAT YOU HAVE IS A RIGHT TRIANGLE WITH A BASE OF 12.27 FT & A SIDE OF 7.5 FEET.
USING THE POINTS (0,0) & (12.27,7.5) WE GET A SLOPE OF:
(12.27-0)/(7.5-0)=12.27/7.5=1.64 FOR THE SLOPE.